OpenStudy (anonymous):

Solve -x + 6 = x - 8 Solve -2(2x - 5) = 2

OpenStudy (anonymous):

isolate x

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

get x on only one side of the equation -x + 6 =x - 8 add x to both sides and add 8 to both sides

OpenStudy (anonymous):

-x+6=x-8 First, move numbers to one side. In this case, subtract 6 from both sides: -x=x-2 Then, move the variables to one side. In this case, subtract x from both sides: -2x=2 After that, divide by -2 on both sides: x= -1

OpenStudy (anonymous):

so it looks like 14 = 2x then divide by 2 u get

OpenStudy (anonymous):

you know...when multiple people post it makes it a lot more confusing for one

OpenStudy (anonymous):

oh okay, what about for the 2nd one?

OpenStudy (anonymous):

i tried and i did it wrong

OpenStudy (anonymous):

well that's distributive

OpenStudy (anonymous):

i kno, i did that and got -4x-5=2

OpenStudy (anonymous):

-4x + 10 = 2 -10 -10 -4x = -8 divide by -4 u get? :3

OpenStudy (anonymous):

ohhh!! wouldnt it be 2?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

- x + 6 = x - 8 -2x = -14 ( -ve's cancel out) 2x = 14 x=7 -2(2x - 5) = 2 -4x + 10 = 2 -4x = 2 - 10 -4x = -8 (-ve's cancel out) 4x=8 x=2 Hope this helps :)

OpenStudy (anonymous):

thanks it does :)